# Proving the existence of localized patterns and saddle node bifurcations in 1D activator-inhibitor type models

### v1

@@ 2.1 Periodic Spaces
In this section, we recall some notations introduced in Section 2.4 of unbounded_domain_cadiot . In particular, the objects defined in the above sections have a corresponding representation in Fourier series when restricted to a bounded domain. Specifically, we define our bounded domain of interest $\Omega_{0}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}(-d,d)$ where $1\leq d<\infty$. Moreover, denote $\tilde{n}=\frac{n}{2d}\in\mathbb{R}$ for all $n\in\mathbb{Z}$. As for the continuous case, we want to restrict to Fourier series representing even functions. Let $\mathbf{u}_{n}=((u_{1})_{n},(u_{2})_{n})$ be the $n$ Fourier coefficient of $\mathbf{u}$. In terms of Fourier coefficients, the even restriction reads
@@
For a bounded operator $K:\ell^{2}_{e}\to\ell^{2}_{e}$ (resp. $K:\ell^{2}(\mathbb{N}_{0})\to\ell^{2}(\mathbb{N}_{0})$), $K^{*}$ denotes the adjoint of $K$ in $\ell^{2}_{e}$ (resp. $\ell^{2}(\mathbb{N}_{0})$). Now, similarly as what is achieved gs_cadiot_blanco , we define $\gamma\penalty 10000\ :\penalty 10000\ L^{2}_{e}(\mathbb{R})\to\ell^{2}(\mathbb{N}_{0})$ as
@@
Then, recall similar notations from unbounded_domain_cadiot
@@
The proof is obtained following similar steps as the ones of Lemma 3.2 in unbounded_domain_cadiot . ∎
@@ 3.1 Radii-Polynomial Theorem
The proof can be found in gs_cadiot_blanco . ∎
@@ 3.2 Construction of $\overline{\mathbf{u}}$
In this section, we discuss the construction of $\overline{\mathbf{u}}$, which is an approximate solution to (22). This is generally a challenging problem. In the specific case of (1), there is a rich variety of work on computing numerical solutions to reaction diffusion systems (e.g alsaadi_spikes ; AlSaadi_unifying_framework ; Champneys_Bistability ). Our approach relies on $\overline{\mathbf{u}}$ being constructed numerically on ${\Omega_{0}}=(-d,d)$ thanks to its Fourier coefficients representation. Fix $N\in\mathbb{N}$ to be the size of our numerical approximation for linear operators (i.e. matrices) and $N_{0}\in\mathbb{N}$ to be the one of our Fourier coefficients approximations (i.e. vectors). Now, given $\mathcal{N}\in\mathbb{N}$, let us introduce the following projection operators
@@
for all $\mathbf{U}=\mathbf{u}_{n\in\mathbb{N}_{0}}\in\ell^{2}_{e}$. To obtain a numerical approximation of a pattern, we looked at the numerical solutions found in AlSaadi_unifying_framework ; Champneys_Bistability and tried to recover them as a Fourier series representation using guesses of the form:
@@
Then we tuned each parameter to replicate the pulse solutions. $\alpha_{i},i=1,2$ controls the stiffness of each spike, and $\beta_{i}$ their amplitude. Once we found a promising candidate, we applied Newton’s method to obtain a better approximate solution of (35). Then we applied the change of variable introduced in Section 1 to get a candidate solution of (2). The multi-pulse solution was obtained through numerical continuation of a candidate in the Brusselator equation, using the AUTO continuation software auto_article . From here, we compute a cosine Fourier sequence approximation of $\mathbf{u}_{0}=(\bar{u}_{1,0},\bar{u}_{2,0})$, $\mathbf{U}_{0}=\gamma(\bar{\mathbf{u}}_{0})\in\ell^{2}_{e}$. We note that $\mathbf{U}_{0}$ satisfies the identity $\mathbf{U}_{0}=\text{{$\Pi$}}^{\leq N_{0}}\mathbf{U}_{0}$ as it is a Fourier sequence with finitely many coefficients.
@@ 3.3 The Operator $\mathbb{A}$
In this section, we focus our attention to the operator $\mathbb{A}:L^{2}_{e}\rightarrow\mathcal{H}_{e}$ required by Theorem 3.1. $\mathbb{A}$ is an approximate inverse of $D\mathbb{f}(\overline{\mathbf{u}})$. We now present its construction following a similar argument as in gs_cadiot_blanco . First we observe that $\mathbb{L}:\mathcal{H}_{e}\rightarrow L^{2}_{e}$ is an isometric isomorphism. Therefore building $\mathbb{A}:L^{2}_{e}\rightarrow\mathcal{H}_{e}$ is equivalent to building $\mathbb{B}:L^{2}_{e}\rightarrow L^{2}_{e}$ approximating the inverse of $D\mathbb{F}(\overline{\mathbf{u}})\mathbb{L}^{-1}$, and setting $\mathbb{A}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\mathbb{L}^{-1}\mathbb{B}$. We will then construct $\mathbb{B}$ using Lemma 2.3 and a bounded linear operator on Fourier coefficients. First we construct $B^{N}$ a numerical approximate inverse to $\text{{$\Pi$}}^{\leq N}Df(\overline{\mathbf{U}})\mathscr{l}^{-1}\text{{$\Pi$}}^{\leq N}$. We choose $B^{N}$ satisfying $B^{N}=\text{{$\Pi$}}^{\leq N}B^{N}\text{{$\Pi$}}^{\leq N}$, hence it admits a matrix representation. We now describe further its construction. Due to the form of $g$, we obtain
@@
The proof can be found in unbounded_domain_cadiot . ∎
@@ 4 Computing the Bounds for Localized Patterns
In this section, we provide explicit formulas for the bounds $\mathcal{Y}_{0},\mathcal{Z}_{1},$ and $\mathcal{Z}_{2}$ in Theorem 3.1. Before we begin the computation of the bounds, we recall some preliminary notations and results from gs_cadiot_blanco . First, given $u\in L^{\infty}(\mathbb{R}^{2})$, denote by
@@
for a given $(x_{1},x_{2},x_{3},x_{4})\in\mathbb{R}^{4}$. For the purpose of our analysis, we recall Lemma 4.2 from gs_cadiot_blanco which allows us to estimate the norm of an operator by block in terms of the norms of its individual blocks.
@@
The proof can be found in gs_cadiot_blanco . ∎
Now, we provide explicit formulas for the bounds of Theorem 3.1 in the following list of lemmas. In particular, each formula relies on finite dimensional computations (that is vector or matrix norms), which can be rigorously evaluated thanks to computer-assisted techniques (cf. julia_blanco_fassler ). This justifies how specific choices for the approximate objects $\bar{\mathbf{u}}$ and $\mathbb{A}$.
@@
We follow the same steps as those of Lemma 4.3 in gs_cadiot_blanco . In particular, we combine (13),the operator $\mathbb{A}$, and the properties of $\gamma$ to get
@@
We then obtain the result using the steps of Lemma 4.11 in unbounded_domain_cadiot . ∎
@@
In fact, notice that $g_{3}(\mathbf{u},\mathbf{u},\mathbf{u})=\mathbb{g}_{3}(\mathbf{u})$. Moreover, we have the symmetry $g_{3}(\mathbf{u},\mathbf{v},\mathbf{w})=g_{3}(\mathbf{v},\mathbf{u},\mathbf{w})$ for all $\mathbf{u},\mathbf{v},\mathbf{w}\in\mathcal{H}_{e}$. Following similar steps to those done in Lemma 4.5 of gs_cadiot_blanco , we obtain
@@
where we also used Lemma 2.2 in the last inequality. Then, using the properties of $\Gamma$, we similar steps to those used in Lemma 4.5 of gs_cadiot_blanco to estimate
@@ 4.1 The Bound $\mathcal{Z}_{1}$
where the final step followed from Theorem 3.5 of unbounded_domain_cadiot and the definition of $\mathbb{B}$. For the final term in (68), we use submultiplicativity to obtain
@@
where the last step followed by Parseval’s Identity and Lemma 4.6 of gs_cadiot_blanco . ∎
@@
The proof follows from Lemma 6.5 of unbounded_domain_cadiot . In particular, each $\mathcal{Z}_{u,k,j}$ can be computed using the results of the aforementioned lemma. ∎
@@ 4.2 Constructive existence proofs of localized solutions
In this section, we apply the Newton-Kantorovich approach presented in Section 3 thanks to the explicit formulas derived in Section 4. In fact, the bounds are evaluated rigorously thanks to the Julia packages IntervalArithmetic.jl julia_interval and RadiiPolynomial.jl julia_olivier . More specifically, the computer-assisted proofs details can be found on julia_blanco_fassler . Now, we present our obtained constructive proofs of existence for localized patterns in (1).
@@
This allows us to compute the $\mathcal{Z}_{2}(r)$ bound defined in Section 4. Finally, using julia_blanco_fassler , we choose $r_{0}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}2\times 10^{-7}$ and define
@@ 5 Saddle-Node Bifurcations : zeros of $\mathbb{F}$
In this manuscript, we present a novel methodology for proving the existence of saddle-node bifurcations on unbounded domains. In fact, our approach relies on the setup described in Section 5 of unbounded_domain_cadiot , which deals with PDEs augmented with extra equations and extra parameters. In fact, in this section we focus on a computer-assisted methodology for proving constructively the existence of non-degenerate zeros of $\mathbb{F}$ given in (7).
@@
As provided in Definition 1.1, non-degenerate zeros of $\mathbb{F}$ provide candidates for a saddle-node bifurcation. Moreover, unbounded_domain_cadiot provides that, given $\tilde{\mathbf{u}}\in\mathcal{H}_{e}$, $D\mathbb{f}(\tilde{\nu},\tilde{\mathbf{u}}):\mathbb{R}\times\mathcal{H}_{e}\to\mathbb{R}\times L^{2}_{e}$ is a Fredholm operator of index zero. The spectral condition of Definition 1.1 will then be verified in the next section. For now, we perform rigorous existence proofs of zeros of (7) in $H_{1}$ using the following theorem.
@@ 5.1 Approximate Inverse
We now briefly recall the setup of Section 5.2 of unbounded_domain_cadiot to build the approximate inverse. Denote $H_{2,{\Omega_{0}}}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\{(\nu,\mathbf{u},\mathbf{w})\in H_{2},\penalty 10000\ \penalty 10000\ \text{supp}(\mathbf{u})\subset\overline{{\Omega_{0}}},\text{supp}(\mathbf{w})\subset\overline{{\Omega_{0}}}\}$ and $H_{1,{\Omega_{0}}}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\{(\nu,\mathbf{u},\mathbf{w})\in H_{1},\penalty 10000\ \penalty 10000\ \text{supp}(\mathbf{u})\subset\overline{{\Omega_{0}}},\text{supp}(\mathbf{w})\subset\overline{{\Omega_{0}}}\}$. By construction, we have $\overline{\mathbf{x}}=(\overline{\nu},\overline{\mathbf{u}},\overline{\mathbf{w}})\in H_{1,{\Omega_{0}}}$. Recall that $\mathcal{B}(Y)$ is the set of bounded linear operators on some space $Y$. Denote by $\mathcal{B}_{\Omega_{0}}(H_{2})$ the following subset of $\mathcal{B}(H_{2})$
@@
As before, we want to approximate the inverse of $D\mathbb{F}(\overline{\mathbf{x}})$ by some operator $\mathbb{A}:H_{2}\to H_{1}$ through the construction of an operator $\mathbb{B}:H_{2}\to H_{2}$, where $\mathbb{A}=\mathbb{L}_{\overline{\nu}}^{-1}\mathbb{B}.$ As described in Section 5 of unbounded_domain_cadiot , we compute a numerical approximate inverse $B^{N}:X_{2}\to X_{2}$ for $\overline{\text{{$\Pi$}}}^{\leq N}DF(\overline{\mathbf{X}})L_{\overline{\nu}}^{-1}\overline{\text{{$\Pi$}}}^{\leq N}$ such that $B^{N}=\overline{\text{{$\Pi$}}}^{\leq N}B^{N}\overline{\text{{$\Pi$}}}^{\leq N}$. Then, we define $\mathbb{A}:H_{2}\to H_{1}$ and $A:X_{2}\to X_{1}$ as

### v2

@@ 2.1 Periodic Spaces
In this section, we recall some notations introduced in Section 2.4 of Cadiot et al. (2024). In particular, the objects defined in the above sections have a corresponding representation in Fourier series when restricted to a bounded domain. Specifically, we define our bounded domain of interest $\Omega_{0}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}(-d,d)$ where $1\leq d<\infty$. Moreover, denote $\tilde{n}=\frac{n}{2d}\in\mathbb{R}$ for all $n\in\mathbb{Z}$. As for the continuous case, we want to restrict to Fourier series representing even functions. Let $\mathbf{u}_{n}=((u_{1})_{n},(u_{2})_{n})$ be the $n$ Fourier coefficient of $\mathbf{u}$. In terms of Fourier coefficients, the even restriction reads
@@
For a bounded operator $K:\ell^{2}_{e}\to\ell^{2}_{e}$ (resp. $K:\ell^{2}(\mathbb{N}_{0})\to\ell^{2}(\mathbb{N}_{0})$), $K^{*}$ denotes the adjoint of $K$ in $\ell^{2}_{e}$ (resp. $\ell^{2}(\mathbb{N}_{0})$). Now, similarly as what is achieved Cadiot and Blanco (2025), we define $\gamma\penalty 10000\ :\penalty 10000\ L^{2}_{e}(\mathbb{R})\to\ell^{2}(\mathbb{N}_{0})$ as
@@
Then, recall similar notations from Cadiot et al. (2024)
@@
The proof is obtained following similar steps as the ones of Lemma 3.2 in Cadiot et al. (2024). ∎
@@ 3.1 Radii-Polynomial Theorem
The proof can be found in Cadiot and Blanco (2025). ∎
@@ 3.2 Construction of $\overline{\mathbf{u}}$
In this section, we discuss the construction of $\overline{\mathbf{u}}$, which is an approximate solution to (22). This is generally a challenging problem. In the specific case of (1), there is a rich variety of work on computing numerical solutions to reaction diffusion systems (e.g Al saadi et al. (2021); Al Saadi et al. (2021); Champneys et al. (2021)). Our approach relies on $\overline{\mathbf{u}}$ being constructed numerically on ${\Omega_{0}}=(-d,d)$ thanks to its Fourier coefficients representation. Fix $N\in\mathbb{N}$ to be the size of our numerical approximation for linear operators (i.e. matrices) and $N_{0}\in\mathbb{N}$ to be the one of our Fourier coefficients approximations (i.e. vectors). Now, given $\mathcal{N}\in\mathbb{N}$, let us introduce the following projection operators
@@
for all $\mathbf{U}=\mathbf{u}_{n\in\mathbb{N}_{0}}\in\ell^{2}_{e}$. To obtain a numerical approximation of a pattern, we looked at the numerical solutions found in Al Saadi et al. (2021); Champneys et al. (2021) and tried to recover them as a Fourier series representation using guesses of the form:
@@
Then we tuned each parameter to replicate the pulse solutions. $\alpha_{i},i=1,2$ controls the stiffness of each spike, and $\beta_{i}$ their amplitude. Once we found a promising candidate, we applied Newton’s method to obtain a better approximate solution of (35). Then we applied the change of variable introduced in Section 1 to get a candidate solution of (2). The multi-pulse solution was obtained through numerical continuation of a candidate in the Brusselator equation, using the AUTO continuation software Doedel et al. (1999). From here, we compute a cosine Fourier sequence approximation of $\mathbf{u}_{0}=(\bar{u}_{1,0},\bar{u}_{2,0})$, $\mathbf{U}_{0}=\gamma(\bar{\mathbf{u}}_{0})\in\ell^{2}_{e}$. We note that $\mathbf{U}_{0}$ satisfies the identity $\mathbf{U}_{0}=\text{{$\Pi$}}^{\leq N_{0}}\mathbf{U}_{0}$ as it is a Fourier sequence with finitely many coefficients.
@@ 3.3 The Operator $\mathbb{A}$
In this section, we focus our attention to the operator $\mathbb{A}:L^{2}_{e}\rightarrow\mathcal{H}_{e}$ required by Theorem 3.1. $\mathbb{A}$ is an approximate inverse of $D\mathbb{f}(\overline{\mathbf{u}})$. We now present its construction following a similar argument as in Cadiot and Blanco (2025). First we observe that $\mathbb{L}:\mathcal{H}_{e}\rightarrow L^{2}_{e}$ is an isometric isomorphism. Therefore building $\mathbb{A}:L^{2}_{e}\rightarrow\mathcal{H}_{e}$ is equivalent to building $\mathbb{B}:L^{2}_{e}\rightarrow L^{2}_{e}$ approximating the inverse of $D\mathbb{F}(\overline{\mathbf{u}})\mathbb{L}^{-1}$, and setting $\mathbb{A}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\mathbb{L}^{-1}\mathbb{B}$. We will then construct $\mathbb{B}$ using Lemma 2.3 and a bounded linear operator on Fourier coefficients. First we construct $B^{N}$ a numerical approximate inverse to $\text{{$\Pi$}}^{\leq N}Df(\overline{\mathbf{U}})\mathscr{l}^{-1}\text{{$\Pi$}}^{\leq N}$. We choose $B^{N}$ satisfying $B^{N}=\text{{$\Pi$}}^{\leq N}B^{N}\text{{$\Pi$}}^{\leq N}$, hence it admits a matrix representation. We now describe further its construction. Due to the form of $g$, we obtain
@@
The proof can be found in Cadiot et al. (2024). ∎
@@ 4 Computing the Bounds for Localized Patterns
In this section, we provide explicit formulas for the bounds $\mathcal{Y}_{0},\mathcal{Z}_{1},$ and $\mathcal{Z}_{2}$ in Theorem 3.1. Before we begin the computation of the bounds, we recall some preliminary notations and results from Cadiot and Blanco (2025). First, given $u\in L^{\infty}(\mathbb{R}^{2})$, denote by
@@
for a given $(x_{1},x_{2},x_{3},x_{4})\in\mathbb{R}^{4}$. For the purpose of our analysis, we recall Lemma 4.2 from Cadiot and Blanco (2025) which allows us to estimate the norm of an operator by block in terms of the norms of its individual blocks.
@@
The proof can be found in Cadiot and Blanco (2025). ∎
Now, we provide explicit formulas for the bounds of Theorem 3.1 in the following list of lemmas. In particular, each formula relies on finite dimensional computations (that is vector or matrix norms), which can be rigorously evaluated thanks to computer-assisted techniques (cf. Blanco et al. (2025)). This justifies how specific choices for the approximate objects $\bar{\mathbf{u}}$ and $\mathbb{A}$.
@@
We follow the same steps as those of Lemma 4.3 in Cadiot and Blanco (2025). In particular, we combine (13),the operator $\mathbb{A}$, and the properties of $\gamma$ to get
@@
We then obtain the result using the steps of Lemma 4.11 in Cadiot et al. (2024). ∎
@@
In fact, notice that $g_{3}(\mathbf{u},\mathbf{u},\mathbf{u})=\mathbb{g}_{3}(\mathbf{u})$. Moreover, we have the symmetry $g_{3}(\mathbf{u},\mathbf{v},\mathbf{w})=g_{3}(\mathbf{v},\mathbf{u},\mathbf{w})$ for all $\mathbf{u},\mathbf{v},\mathbf{w}\in\mathcal{H}_{e}$. Following similar steps to those done in Lemma 4.5 of Cadiot and Blanco (2025), we obtain
@@
where we also used Lemma 2.2 in the last inequality. Then, using the properties of $\Gamma$, we similar steps to those used in Lemma 4.5 of Cadiot and Blanco (2025) to estimate
@@ 4.1 The Bound $\mathcal{Z}_{1}$
where the final step followed from Theorem 3.5 of Cadiot et al. (2024) and the definition of $\mathbb{B}$. For the final term in (68), we use submultiplicativity to obtain
@@
where the last step followed by Parseval’s Identity and Lemma 4.6 of Cadiot and Blanco (2025). ∎
@@
The proof follows from Lemma 6.5 of Cadiot et al. (2024). In particular, each $\mathcal{Z}_{u,k,j}$ can be computed using the results of the aforementioned lemma. ∎
@@ 4.2 Constructive existence proofs of localized solutions
In this section, we apply the Newton-Kantorovich approach presented in Section 3 thanks to the explicit formulas derived in Section 4. In fact, the bounds are evaluated rigorously thanks to the Julia packages IntervalArithmetic.jl Benet and Sanders (2022) and RadiiPolynomial.jl Hénot (2022). More specifically, the computer-assisted proofs details can be found on Blanco et al. (2025). Now, we present our obtained constructive proofs of existence for localized patterns in (1).
@@
This allows us to compute the $\mathcal{Z}_{2}(r)$ bound defined in Section 4. Finally, using Blanco et al. (2025), we choose $r_{0}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}2\times 10^{-7}$ and define
@@ 5 Saddle-Node Bifurcations : zeros of $\mathbb{F}$
In this manuscript, we present a novel methodology for proving the existence of saddle-node bifurcations on unbounded domains. In fact, our approach relies on the setup described in Section 5 of Cadiot et al. (2024), which deals with PDEs augmented with extra equations and extra parameters. In fact, in this section we focus on a computer-assisted methodology for proving constructively the existence of non-degenerate zeros of $\mathbb{F}$ given in (7).
@@
As provided in Definition 1.1, non-degenerate zeros of $\mathbb{F}$ provide candidates for a saddle-node bifurcation. Moreover, Cadiot et al. (2024) provides that, given $\tilde{\mathbf{u}}\in\mathcal{H}_{e}$, $D\mathbb{f}(\tilde{\nu},\tilde{\mathbf{u}}):\mathbb{R}\times\mathcal{H}_{e}\to\mathbb{R}\times L^{2}_{e}$ is a Fredholm operator of index zero. The spectral condition of Definition 1.1 will then be verified in the next section. For now, we perform rigorous existence proofs of zeros of (7) in $H_{1}$ using the following theorem.
@@ 5.1 Approximate Inverse
We now briefly recall the setup of Section 5.2 of Cadiot et al. (2024) to build the approximate inverse. Denote $H_{2,{\Omega_{0}}}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\{(\nu,\mathbf{u},\mathbf{w})\in H_{2},\penalty 10000\ \penalty 10000\ \text{supp}(\mathbf{u})\subset\overline{{\Omega_{0}}},\text{supp}(\mathbf{w})\subset\overline{{\Omega_{0}}}\}$ and $H_{1,{\Omega_{0}}}\stackrel{{\scriptstyle\mbox{\tiny{\raisebox{0.0pt}[0.0pt][0.0pt]{def}}}}}{{=}}\{(\nu,\mathbf{u},\mathbf{w})\in H_{1},\penalty 10000\ \penalty 10000\ \text{supp}(\mathbf{u})\subset\overline{{\Omega_{0}}},\text{supp}(\mathbf{w})\subset\overline{{\Omega_{0}}}\}$. By construction, we have $\overline{\mathbf{x}}=(\overline{\nu},\overline{\mathbf{u}},\overline{\mathbf{w}})\in H_{1,{\Omega_{0}}}$. Recall that $\mathcal{B}(Y)$ is the set of bounded linear operators on some space $Y$. Denote by $\mathcal{B}_{\Omega_{0}}(H_{2})$ the following subset of $\mathcal{B}(H_{2})$
@@
As before, we want to approximate the inverse of $D\mathbb{F}(\overline{\mathbf{x}})$ by some operator $\mathbb{A}:H_{2}\to H_{1}$ through the construction of an operator $\mathbb{B}:H_{2}\to H_{2}$, where $\mathbb{A}=\mathbb{L}_{\overline{\nu}}^{-1}\mathbb{B}.$ As described in Section 5 of Cadiot et al. (2024), we compute a numerical approximate inverse $B^{N}:X_{2}\to X_{2}$ for $\overline{\text{{$\Pi$}}}^{\leq N}DF(\overline{\mathbf{X}})L_{\overline{\nu}}^{-1}\overline{\text{{$\Pi$}}}^{\leq N}$ such that $B^{N}=\overline{\text{{$\Pi$}}}^{\leq N}B^{N}\overline{\text{{$\Pi$}}}^{\leq N}$. Then, we define $\mathbb{A}:H_{2}\to H_{1}$ and $A:X_{2}\to X_{1}$ as